# The CFT of SLE loop measures and the Kontsevich–Suhov conjecture

### v1

@@ 1.1. Background and overview
In this paper, we consider the loop version of SLE [KW16, Zha21] (see also [Wer08, BD16] for the special cases $\kappa=\frac{8}{3},2$), and introduce the main ingredients that will turn the SLE loop measure into a full-fledged CFT. In a nutshell, we construct a structure $(L^{2}(\nu),(\mathbf{L}_{n},\bar{\mathbf{L}}_{n})_{n\in\mathbb{Z}},\mathcal{Q})$, where $\nu$ is the SLEκ loop measure ($\kappa\in(0,4]$), $(\mathbf{L}_{n},\bar{\mathbf{L}}_{n})_{n\in\mathbb{Z}}$ are two commuting representations of the Virasoro algebra with central charge $c_{\mathrm{m}}=1-6(\frac{2}{\sqrt{\kappa}}-\frac{\sqrt{\kappa}}{2})^{2}\leq 1$ acting as (unbounded) operators on $L^{2}(\nu)$, and $\mathcal{Q}$ is a Hermitian form on $L^{2}(\nu)$ such that $\mathbf{L}_{-n}$ is the $\mathcal{Q}$-adjoint of $\mathbf{L}_{n}$ for all $n\in\mathbb{Z}$. In representation theoretic terms, $\mathcal{Q}$ is called the Shapovalov form. It turns out to have a very natural geometric interpretation, consisting in “exchanging” the interior and the exterior of the curve. Our representations are constructed as follows: the Witt algebra defines invariant vector fields on the space of Jordan curves, and the operators are Lie derivatives in these directions. Representations of this flavour (but with substantial differences described in detail in Section 1.3) have been considered in [KY87, KY88, BB03b, Kyt07], seeking an analogue of the Borel-Weil theorem for orbits of the group of diffeomorphisms of the circle. Much closer to our approach are the representations considered in [CP14, GQW24], inspired by the variational formulas of [DS62].
@@
One of the most important aspects of our work is that the Shapovalov form is distinct from the $L^{2}(\nu)$-inner product, which is a symptom of the non-unitarity of the theory. In fact, it is a standard result from the representation theory of the Virasoro algebra that the Shapovalov form is not positive definite for $c_{\mathrm{m}}\leq 1$ (except for a discrete set of values). In physics, non-unitary theories and their relations to loop models have been studied quite intensively over the past few years, mostly by algebraic methods and consistency conditions (crossing symmetry, conformal bootstrap, null-vector equations…) [RS15, PRS16, HJS20, JRS22]. However, contrary to the regime of central charge $\geq 25$, there seems to be a zoo of possible theories when the central charge is less than 1. This underlines the importance of studying concrete representations like we do in this paper. In the probabilistic literature, a closely related model has been studied with a different point of view [AS21, ARS22], namely using a coupling of SLE with Liouville CFT [DKRV16, KRV20, GKRV23] inspired by Sheffield’s quantum zipper [She16] (see also the comprehensive book [BP24] for a detailed and gentle introduction to this difficult topic). In fact, the generating function of the “electrical thickness” computed in [AS21] is the partition function of our Shapovalov form. This type of coupling also appears in theoretical physics: [Ebe23, (2.20)] proposes an inner product on the space of Virasoro conformal blocks at central charge $c_{\mathrm{L}}\geq 25$ by reweighting by the partition function of a CFT with central charge $c_{\mathrm{m}}=26-c_{\mathrm{L}}$. The putative CFT behind this partition function is not clearly identified there, but is conjectured to exist for all values of the central charge $c_{\mathrm{m}}\leq 1$, with a continuous spectrum of representations [Ebe23, Conjecture 1]. These two properties are compatible with our results, and make SLE a natural candidate. The inner product of [Ebe23] can also be interpreted in light of the Polyakov path integral [Pol81], where the matter field is an SLEκ-type measure on pairs of pants decomposition, and the ghosts are the trivial Beltrami differentials. We do not make any precise statement at this stage, but view this as an interesting direction of research.
@@
Acknowledgments We are particularly grateful to E. Peltola for numerous discussions on SLE and a detailed feedback on this manuscript. We thank E. Peltola and B. Wu for bringing the question of uniqueness of MKS measures to our attention and W. Werner for pointing out some references to us. We also thank V. Vargas for discussions on SLE and the conformal bootstrap. Some of this work has been achieved during the fantastic conference “Probability in CFT” organised by J. Aru, C. Guillarmou and R. Rhodes, at the Bernoulli Center, EPFL, whose hospitality is greatly acknowledged. AJ is supported by Eccellenza grant 194648 of the Swiss National Science Foundation and is a member of NCCR SwissMAP.
During the final stages of preparation of the current article, we learnt that M. Gordina, W. Qian and Y. Wang were independently working on a related article [GQW24]. Specifically, they construct representations of the Witt and Virasoro algebras on $L^{2}(\nu)$ which would roughly correspond to our Theorems 1.1 and 1.2 below. They also derive an integration by parts formula and construct the Shapovalov form, which would correspond to Theorem 1.3 below. Their approach partly differs from ours, and we invite the reader to check it out in details. We thank them for bringing the references [DS62, CP14] to our knowledge.
@@ 1.7. Uniqueness of restriction measures
For any domain $D\subset\widehat{\mathbb{C}}$, we denote by $\mu^{\mathrm{loop}}_{D}$ the Brownian loop measure in $D$, which is a measure on Brownian-type loops introduced by Lawler–Werner [LW04]; see below Proposition 4.5 for details. For any Borel sets $A,B\subset D$, we will denote by $\Lambda_{D}(A,B)$ the $\mu^{\mathrm{loop}}_{D}$-mass of loops touching both $A$ and $B$. Alternatively, and to make the link with [KS07], $\mathrm{exp}(\Lambda_{D}(A,B))$ can be defined as a ratio of $\zeta$-regularised determinants; see [Dub09, Proposition 2.1]. When $\partial D$ is non polar and $A$ and $B$ are disjoint closed sets, $\Lambda_{D}(A,B)$ is a finite quantity. If $D=\widehat{\mathbb{C}}$ and $A$ and $B$ are disjoint non polar subsets, this quantity is infinite but can be renormalised as follows (see [FL13, Theorem 1.1]):
@@
Let $\nu$ be a Borel measure on the space $\mathcal{J}$ of Jordan curves in the Riemann sphere, and denote by $\nu|_{\mathcal{J}_{D}}$ its restriction to the set $\mathcal{J}_{D}$ of loops in $D$, for any domain $D$. We say that
  1. (1)
    $\nu$ is conformally invariant if $\Phi^{*}\nu=\nu$ for all Möbius transformation $\Phi$ of the Riemann sphere.
  2. (2)
@@
    $\nu$ satisfies conformal restriction (with central charge $c_{\mathrm{m}}$) if for all simply connected domains $D_{1},D_{2}\subset\widehat{\mathbb{C}}$ with $\bar{D}_{2}\subset D_{1}$, $\nu|_{\mathcal{J}_{D_{2}}}$ is absolutely continuous with respect to $\nu|_{\mathcal{J}_{D_{1}}}$ and
    (1.21) $\frac{\mathrm{d}\nu|_{\mathcal{J}_{D_{2}}}}{\mathrm{d}\nu|_{\mathcal{J}_{D_{1}}}}=\mathds{1}_{\eta\subset D_{2}}\mathrm{exp}\left(\frac{c_{\mathrm{m}}}{2}\left(\Lambda^{*}(\eta,D_{1}^{c})-\Lambda^{*}(\eta,D_{2}^{c})+\Lambda_{D_{1}}(\eta,D_{2}^{c})\right)\right).$
@@
Measures satisfying these properties have a long history. In [AM01], Airault and Malliavin introduced the notion of unitarising measures for the Virasoro algebra from an axiomatic point of view [^2] In our notations, their definition would require $\mathbf{L}_{n}^{*}=\mathbf{L}_{-n}$ on the $L^{2}$-space of the measure, which is impossible due to the lack of unitarity when the central charge is less than 1. The correct notion of unitarising measure is the one studied in this paper where we construct the Shapovalov form as a Hermitian form distinct from the $L^{2}$-product., leaving the construction of such measures open. Interestingly, this notion appeared almost simultaneously as Schramm’s original paper on SLE [Sch00]. In parallel, Lawler, Schramm and Werner showed the restriction property for chordal SLE [LSW03]. Then, Kontsevich and Suhov related the restriction property to unitarising measures and conjectured that, for each central charge $c_{\mathrm{m}}\leq 1$, there exists a unique (up to multiplicative constant) measure on Jordan curves satisfying the properties of Definition 1.8 [KS07, Conjecture 1]. The construction of such loop measures was first done in the special cases $c_{\mathrm{m}}=0$ [Wer08] and $c_{\mathrm{m}}=-2$ [BD16]. A construction for the range $c_{\mathrm{m}}\in(0,1]$ was later provided by Kemppainen and Werner [KW16] based on Conformal Loop Ensemble (counting measure of nested CLE). Finally, Zhan’s construction [Zha21] of the SLE loop measure covers the existence part of Kontsevich–Suhov’s conjecture in the whole region $c_{\mathrm{m}}\leq 1$. In the range $c_{\mathrm{m}}\in(0,1]$, Zhan’s construction is closely related to the argument of Kemppainen–Werner showing reversibility [KW16], and indeed the two measures differ only by a multiplicative constant [AS21, Theorem 2.18]. We also mention that the articles [Ben16, MP24] study more general measures where the conformal restriction anomaly in (1.21) is not necessarily given by the Brownian loop measure. Here, we show the uniqueness part for loops in the sphere.
Theorem 1.9.
@@
Let $c_{\mathrm{m}}\leq 1$. Up to a positive multiplicative constant, there exists a unique locally finite, non-zero Borel measure on the space of Jordan curves in the Riemann sphere satisfying conformal invariance and conformal restriction with central charge $c_{\mathrm{m}}$. This measure is the SLEκ loop measure for the unique $\kappa\in(0,4]$ such that $c_{\mathrm{m}}=1-6(\frac{2}{\sqrt{\kappa}}-\frac{\sqrt{\kappa}}{2})^{2}$.
@@
This characterisation of SLE in terms of an infinitesimal restriction property could offer a new perspective on cutting/welding random surfaces. We plan to investigate this in a future work and explain this in more details in Section 7.7.
@@ 2.3. Witt algebra
In this section, we construct and study two commuting representations $(\mathcal{L}_{\mathrm{v}})_{\mathrm{v}\in\mathbb{C}(z)\partial_{z}}$ and (¯$(\bar{\mathcal{L}}_{\mathrm{v}})_{\mathrm{v}\in\mathbb{C}(z)\partial_{z}}$ of the Witt algebra as endomorphisms of the dense subspace $\mathcal{C}$ (2.5) of $L^{2}(\nu)$. The results of this section are similar to [GQW24, Section 3.2] and [CP14, Section 3], which are obtained from the variational formulas of [DS62]. Since we were not aware of this reference, we rederived these formulas from scratch using a different method, namely Ahlfors’ variational formulas for quasiconformal mappings [Ahl87].
@@ 4.2. Integration by parts for the SLE loop measure in the unit disc
Here, “hit both $\text{hit both }\eta\text{ and }\mathbb{D}\setminus D$” is the event that the sample of $\mu^{\mathrm{loop}}_{D}$ intersects both the trace of $\eta$ and the boundary of $\mathbb{D}\setminus D$. This transparent notation will be used throughout the proof.
@@
In Lemmas 4.6 and 4.7 below, we will show that
@@
The rest of the section is dedicated to the proof of (4.9). Up to considering the vector field $e^{i\theta}\mathrm{v}$ for $|\theta|<\delta$, it suffices to consider the case of small positive values of $t$. We decompose the proof into two steps, corresponding to Lemmas 4.6 and 4.7 respectively.
@@
We now relate $\mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit }\eta)$ to the quantity appearing in the statement of Proposition 4.5. Recall that $q(\eta)$ and $\psi_{\eta}$ are introduced in Notation 4.4 and that $U(q)$ is defined in (4.6).
@@
Combining Lemmas 4.6 and 4.7, we obtain

### v2

@@ 1.1. Background and overview
In this paper, we consider the loop version of SLE [KW16, Zha21] (see also [Wer08, BD16] for the special cases $\kappa=\frac{8}{3},2$), and introduce the main ingredients that will turn the SLE loop measure into a full-fledged CFT. In a nutshell, we construct a structure $(L^{2}(\nu),(\mathbf{L}_{n},\bar{\mathbf{L}}_{n})_{n\in\mathbb{Z}},\mathcal{Q})$, where $\nu$ is the SLEκ loop measure ($\kappa\in(0,4]$), $(\mathbf{L}_{n},\bar{\mathbf{L}}_{n})_{n\in\mathbb{Z}}$ are two commuting representations of the Virasoro algebra with central charge $c_{\mathrm{m}}=1-6(\frac{2}{\sqrt{\kappa}}-\frac{\sqrt{\kappa}}{2})^{2}\leq 1$ acting as (unbounded) operators on $L^{2}(\nu)$, and $\mathcal{Q}$ is a Hermitian form on $L^{2}(\nu)$ such that $\mathbf{L}_{-n}$ is the $\mathcal{Q}$-adjoint of $\mathbf{L}_{n}$ for all $n\in\mathbb{Z}$. In representation theoretic terms, $\mathcal{Q}$ is called the Shapovalov form. It turns out to have a very natural geometric interpretation, consisting in “exchanging” the interior and the exterior of the curve. Our representations are constructed as follows: the Witt algebra defines invariant vector fields on the space of Jordan curves, and the operators are Lie derivatives in these directions. Representations of this flavour (but with substantial differences described in detail in Section 1.3) have been considered in [KY87, KY88, BB03b, Kyt07], seeking an analogue of the Borel-Weil theorem for orbits of the group of diffeomorphisms of the circle. Much closer to our approach are the representations considered in [CP14, GQW24], inspired by the variational formulas of [DS63].
@@
One of the most important aspects of our work is that the Shapovalov form is distinct from the $L^{2}(\nu)$-inner product, which is a symptom of the non-unitarity of the theory. In fact, it is a standard result from the representation theory of the Virasoro algebra that the Shapovalov form is not positive definite for $c_{\mathrm{m}}\leq 1$ (except for a discrete set of values). In physics, non-unitary theories and their relations to loop models have been studied quite intensively over the past few years, mostly by algebraic methods and consistency conditions (crossing symmetry, conformal bootstrap, null-vector equations…) [RS15, PRS16, HJS20, JRS23]. However, contrary to the regime of central charge $\geq 25$, there seems to be a zoo of possible theories when the central charge is less than 1. This underlines the importance of studying concrete representations like we do in this paper. In the probabilistic literature, a closely related model has been studied with a different point of view [AS21, ARS22], namely using a coupling of SLE with Liouville CFT [DKRV16, KRV20, GKRV23] inspired by Sheffield’s quantum zipper [She16] (see also the comprehensive book [BP24] for a detailed and gentle introduction to this difficult topic). In fact, the generating function of the “electrical thickness” computed in [AS21] is the partition function of our Shapovalov form. This type of coupling also appears in theoretical physics: [Ebe23, (2.20)] proposes an inner product on the space of Virasoro conformal blocks at central charge $c_{\mathrm{L}}\geq 25$ by reweighting by the partition function of a CFT with central charge $c_{\mathrm{m}}=26-c_{\mathrm{L}}$. The putative CFT behind this partition function is not clearly identified there, but is conjectured to exist for all values of the central charge $c_{\mathrm{m}}\leq 1$, with a continuous spectrum of representations [Ebe23, Conjecture 1]. These two properties are compatible with our results, and make SLE a natural candidate. The inner product of [Ebe23] can also be interpreted in light of the Polyakov path integral [Pol81], where the matter field is an SLEκ-type measure on pairs of pants decomposition, and the ghosts are the trivial Beltrami differentials. We do not make any precise statement at this stage, but view this as an interesting direction of research.
@@
Acknowledgments We are particularly grateful to E. Peltola for numerous discussions on SLE and a detailed feedback on this manuscript. We thank E. Peltola and B. Wu for bringing the question of uniqueness of MKS measures to our attention, and W. Werner for pointing out the relevance of our results to the reversibility and the duality of SLE. We also thank V. Vargas for discussions on SLE and the conformal bootstrap. Some of this work has been achieved during the fantastic conference “Probability in CFT” organised by J. Aru, C. Guillarmou and R. Rhodes, at the Bernoulli Center, EPFL, whose hospitality is greatly acknowledged. AJ is supported by Eccellenza grant 194648 of the Swiss National Science Foundation and is a member of NCCR SwissMAP.
During the final stages of preparation of the current article, we learnt that M. Gordina, W. Qian and Y. Wang were independently working on a related article [GQW24]. Specifically, they construct representations of the Witt and Virasoro algebras on $L^{2}(\nu)$ which would roughly correspond to our Theorems 1.1 and 1.2 below. They also derive an integration by parts formula and construct the Shapovalov form, which would correspond to Theorem 1.3 below. Their approach partly differs from ours, and we invite the reader to check it out in details. We thank them for bringing the references [DS63, CP14] to our knowledge.
@@ 1.7. Uniqueness of restriction measures
For any domain $D\subset\widehat{\mathbb{C}}$, we denote by $\mu^{\mathrm{loop}}_{D}$ the Brownian loop measure in $D$, which is a measure on Brownian-type loops introduced by Lawler–Werner [LW04]; see below Proposition 4.6 for details. For any Borel sets $A,B\subset D$, we will denote by $\Lambda_{D}(A,B)$ the $\mu^{\mathrm{loop}}_{D}$-mass of loops touching both $A$ and $B$. Alternatively, and to make the link with [KS07], $\mathrm{exp}(\Lambda_{D}(A,B))$ can be defined as a ratio of $\zeta$-regularised determinants; see [Dub09b, Proposition 2.1]. When $\partial D$ is non polar and $A$ and $B$ are disjoint closed sets, $\Lambda_{D}(A,B)$ is a finite quantity. If $D=\widehat{\mathbb{C}}$ and $A$ and $B$ are disjoint non polar subsets, this quantity is infinite but can be renormalised as follows (see [FL13, Theorem 1.1]):
@@
Fix $c_{\mathrm{m}}\leq 1$. Let $\nu$ be a Borel measure on the space $\mathcal{J}$ (resp. $\mathcal{J}_{0,\infty}$) of Jordan curves in the Riemann sphere (resp. separating 00 from $\infty$). For any simply connected domain $D$, denote by
(1.21) $\mathrm{d}\nu_{D}(\eta):=\mathds{1}_{\eta\subset D}e^{\frac{c_{\mathrm{m}}}{2}\Lambda^{*}(\eta,\partial D)}\mathrm{d}\nu(\eta).$
@@
We say that $\nu$ is a restriction measure on $\mathcal{J}$ (resp. a weak restriction measure on $\mathcal{J}_{0,\infty}$), with central charge $c_{\mathrm{m}}$, if for any simply connected domains $D_{1},D_{2}$ (resp. both containing 00 and included in $^{}^\widehat{\mathbb{C}}\setminus\{\infty\}$) and any conformal transformation $\Phi:D_{1}\to D_{2}$, $\nu_{D_{2}}=\Phi^{*}\nu_{D_{1}}$.
Note that, in the above definition of weak restriction measures on $\mathcal{J}_{0,\infty}$, there is no assumption of invariance under conformal maps which exchange the roles of 00 and $\infty$. The word “weak” refers to this weak notion of conformal invariance. This will be used in Section 7.1 in order to give a new proof of reversibility of SLE.
@@
Restriction measures on $\mathcal{J}$, also referred to as Malliavin–Kontsevich–Suhov measures (or simply MKS measures), have a long history. In [AM01], Airault and Malliavin introduced the notion of unitarising measures for the Virasoro algebra from an axiomatic point of view [^2] In our notations, their definition would require $\mathbf{L}_{n}^{*}=\mathbf{L}_{-n}$ on the $L^{2}$-space of the measure, which is impossible due to the lack of unitarity when the central charge is less than 1. The correct notion of unitarising measure is the one studied in this paper where we construct the Shapovalov form as a Hermitian form distinct from the $L^{2}$-product., leaving the construction of such measures open. Interestingly, this notion appeared almost simultaneously as Schramm’s original paper on SLE [Sch00]. In parallel, Lawler, Schramm and Werner showed the restriction property for chordal SLE [LSW03]. Then, Kontsevich and Suhov related the restriction property to unitarising measures and conjectured that, for each central charge $c_{\mathrm{m}}\leq 1$, there exists a unique (up to multiplicative constant) measure on Jordan curves satisfying the properties of Definition 1.8 [KS07, Conjecture 1]. The construction of such loop measures was first done in the special cases $c_{\mathrm{m}}=0$ [Wer08] and $c_{\mathrm{m}}=-2$ [BD16]. A construction for the range $c_{\mathrm{m}}\in(0,1]$ was later provided by Kemppainen and Werner [KW16] based on Conformal Loop Ensemble (counting measure of nested CLE). Finally, Zhan’s construction [Zha21] of the SLE loop measure covers the existence part of Kontsevich–Suhov’s conjecture in the whole region $c_{\mathrm{m}}\leq 1$. In the range $c_{\mathrm{m}}\in(0,1]$, Zhan’s construction is closely related to the argument of Kemppainen–Werner showing reversibility [KW16], and indeed the two measures differ only by a multiplicative constant [AS21, Theorem 2.18]. We also mention that the articles [Ben16, MP24] study more general measures where the conformal restriction anomaly in (1.21) is not necessarily given by the Brownian loop measure. Here, we show the uniqueness part for loops in the sphere.
Theorem 1.9 (Uniqueness of restriction measures).
@@
Let $c_{\mathrm{m}}\leq 1$. Up to a positive multiplicative constant, there exists a unique restriction measure on $\mathcal{J}$ (resp. weak restriction measure on $\mathcal{J}_{0,\infty}$), with central charge $c_{\mathrm{m}}$, that is non-zero and locally finite. This measure is the SLEκ loop measure (resp. the SLEκ loop measure restricted to the space $\mathcal{J}_{0,\infty}$), for the unique $\kappa\in(0,4]$ such that $c_{\mathrm{m}}=1-6(\frac{2}{\sqrt{\kappa}}-\frac{\sqrt{\kappa}}{2})^{2}$.
@@
This characterisation of SLE in terms of an infinitesimal restriction property could offer a new perspective on cutting/welding random surfaces. We plan to investigate this in a future work and explain this in more details in Section 8.7. We also discuss two consequences of this uniqueness result in Section 7; namely, the reversibility and the duality of SLE loop measures.
@@ 1.8. Organisation and notations
  • $\bullet$
    Section 7: we discuss two consequences of our uniqueness result: the reversibility and the duality of SLE loop measures.
@@ 2.3. Witt algebra
In this section, we construct and study two commuting representations $(\mathcal{L}_{\mathrm{v}})_{\mathrm{v}\in\mathbb{C}(z)\partial_{z}}$ and (¯$(\bar{\mathcal{L}}_{\mathrm{v}})_{\mathrm{v}\in\mathbb{C}(z)\partial_{z}}$ of the Witt algebra as endomorphisms of the dense subspace $\mathcal{C}$ (2.5) of $L^{2}(\nu)$. The results of this section are similar to [GQW24, Section 3.2] and [CP14, Section 3], which are obtained from the variational formulas of [DS63]. Since we were not aware of this reference, we rederived these formulas from scratch using a different method, namely Ahlfors’ variational formulas for quasiconformal mappings [Ahl87].
@@ 4. Integration by parts
Remark 4.3.
In Section 3, we defined the modules $\mathcal{W}_{\lambda}$ as algebraic subspaces of $\mathbb{C}[(a_{m},\bar{a}_{m})_{m\geq 1}]$, and the representations $(\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}_{n}^{\lambda})_{n\in\mathbb{Z}}$ as endomorphisms acting on these spaces. Such definitions are sufficient for the purpose of characterising these modules algebraically. Now, we can also consider the closure of $\mathcal{W}_{\lambda}$ with respect to the $L^{2}(\nu^{\#})$; this is a closed subspace of $L^{2}(\nu^{\#})$, which is proper if and only if $\lambda$ lies on the Kac table. The operators $\mathbf{L}_{n}^{\lambda},\bar{\mathbf{L}}_{n}^{\lambda}$ extend to unbounded operators on the closure, and one can deduce from the above integration by parts formula that they are closable (the adjoints are densely defined).
@@ 4.2. Integration by parts for the SLE loop measure in the unit disc
Here, “hit both $\text{hit both }\eta\text{ and }\mathbb{D}\setminus D$” is the event that the sample of $\mu^{\mathrm{loop}}_{D}$ intersects both the trace of $\eta$ and the boundary of $\mathbb{D}\setminus D$. This transparent notation will be used throughout the proof. In Definition 1.8, the conformal restriction property of the SLE loop measure was phrased slightly differently since we started with a measure on the sphere. These two restriction properties agree since for any simply connected domains $D_{1},D_{2}$ with $\overline{D_{2}}\subset D_{1}$ and for any Jordan curve $\eta$ in $D_{2}$,
(4.8) $\Lambda_{D_{1}}(\eta,D_{2}^{c})=\Lambda^{*}(\eta,D_{2}^{c})-\Lambda^{*}(\eta,D_{1}^{c}).$
This last relation in turn simply follows from the restriction property of the Brownian loop measure.
@@
In Lemmas 4.7 and 4.8 below, we will show that
@@
The rest of the section is dedicated to the proof of (4.10). Up to considering the vector field $e^{i\theta}\mathrm{v}$ for $|\theta|<\delta$, it suffices to consider the case of small positive values of $t$. We decompose the proof into two steps, corresponding to Lemmas 4.7 and 4.8 respectively.
@@
We now relate $\mu^{\mathrm{bub}}_{\mathbb{D},e^{i\theta}}(\text{hit }\eta)$ to the quantity appearing in the statement of Proposition 4.6. Recall that $q(\eta)$ and $\psi_{\eta}$ are introduced in Notation 4.5 and that $U(q)$ is defined in (4.6).
@@
Combining Lemmas 4.7 and 4.8, we obtain